OpenStudy (anonymous):

what velocity in m/s did an airplane take off from a landing field that has a run of 500m. if it stats from rest, moves with constant acceleration and makes the run in 30 sec?

OpenStudy (jfraser):

use the equations of motion (the terrific trio)\[\Delta x = V_0*t + \frac{1}{2} a*t^2\]\[V_f^2 = V_0^2 + 2a*\Delta x\]and\or \[V_f = V_0 + a*t\]each equation uses the same set of pieces, use algebra to find the missing ones.

OpenStudy (anonymous):

question.. so im going to substitute the Δx to Vf^2=V0^2+2a∗Δx??

OpenStudy (anonymous):

or from Δx i should get acceleration then use Vf=V0+a∗t to get vf?

OpenStudy (jfraser):

it actually doesn't matter which one you use, because some equations will use some pieces, and other equations will use other pieces. you know a Dx, t, and a V0, so pick the equation that most fits

OpenStudy (jfraser):

and you might have to use 2

OpenStudy (anonymous):

oh.. ok ok acceleration is the only missing piece right?

OpenStudy (jfraser):

it's one of the missing pieces, Vf is the other